OpenStudy (anonymous):

A shopkeeper buys a number of books for Rs.1200. If he had bought 10 more books for the same amount, each book would have cost Rs. 20 less. How many books did he buy.

OpenStudy (anonymous):

Let no. books be x. total cost of books = 1200 so, cost of 1 book = (1200 / x)

OpenStudy (anonymous):

case 2 - total cost of books = 1200 no. of books = ( x + 10 ) so, cost of 1 book = ( 1200 / x +10 )

OpenStudy (anonymous):

given, original cost - 20 = new cost so, ( 1200 / x ) - 20 = (1200 / x +10) on solving, eq. formed - x^2 + 10 x - 600 = 0 by, quadratic formula, x = -30 or 20 as no. of books cannot be negative, so , x cannot be equal to -30 so, x = 20 books

OpenStudy (anonymous):

let x be number of books bought so cost of one book= 1200/x in other case cost of one book = 1200/x+10 thus (1200/x)-(1200/x+10)=20 quadratic equation formed